# AnyV2V: A Tuning-Free Framework For Any Video-to-Video Editing Tasks

### v3

@@ 4.2 Evaluation Results
As shown in Figure 3 and 4, AnyV2V can perform the following video editing tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation. we compare AnyV2V against three baseline models Tune-A-Video [53], TokenFlow [16] and FLATTEN [12] for (1) prompt-based editing. Since there exists no publicly available baseline method for task (2) (3) (4), we evaluate the performance of three I2V generation models under AnyV2V. We included a more comprehensive evaluation in Appendix B.
@@
Prompt-based Editing Unlike the baseline methods [53, 16, 12] that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged. Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix [4], AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue. Additionally, it maintains the original video’s background and fidelity, whereas, in comparison, baseline methods often alter the color tone and shape of objects, as illustrated in Figure 3. Also, for motion tasks such as adding snowing weather, I2V models from AnyV2V provide inherent support for animating the snowing while the baseline methods would result in flickering. For quantitative evaluations, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting, and also compute the metrics CLIP-Text for text alignment and CLIP-Image for temporal consistency. In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model [41] and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames. Table 2 shows that AnyV2V generally achieves high text-alignment and temporal consistency, while AnyV2V with I2VGen-XL backbone is the most preferred method because it does not edit the video precisely.
@@
Style Transfer, Subject-Driven Editing and Identity Manipulation For these novel tasks, we stress the alignment with the reference image instead of the text prompt. As shown in Figure 4, we can observe that for task (2), AnyV2V can capture one particular style that is tailor-made, and even if the style is not learned by the text encoder. In the examples, AnyV2V captures the style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately. For task (3), AnyV2V can replace the subject in the video with other subjects even if the new subject differs from the original subject. In the examples, a cat is replaced with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. A car is replaced by our desired car while the wheel is still spinning in the edited video. For task (4), AnyV2V can swap a person’s identity to anyone. We report both human evaluation results and find that AnyV2V with I2VGen-XL backbone is the most preferred method in terms of reference alignment and overall performance, plotted in Table 5, which is in Appendix B.
@@
I2V Backbones We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively. It has a good generalization ability to produce consistent motions in the video with high visual quality. AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its training data, thus harming the visual quality. AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models [10, 42, 59] are mostly trained on video data that contains only 16 frames. To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V [42] and SEINE [10]. However, we find that such an experiment setup cannot maintain semantic consistency in our case. As many works in extending video generation exploit the initial latent to generate longer video [40, 54], we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output. Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
@@
Figure 5: AnyV2V can edit video length beyond the training frame while maintaining motion consistency. The first row is the source video frames while the second rows are the edited. The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix [4].
@@ A.1 U-Net Layers for Feature Injection
To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL [59], ConsistI2V [42] and SEINE [10]. Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e. average attention weights for all other tokens attending to the current token). The results are shown in Figure 7.
@@ A.2 Ablation Analysis on Feature Injection Thresholds
Appendix B Evaluation
B.1 Tasks Definition
  • 1.
  • 2.
  • 3.
  • 4.
B.2 Quantitative Evaluations
@@
For (1) prompt-based editing, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting. We compare AnyV2V against three baseline models: Tune-A-Video [53], TokenFlow [16] and FLATTEN [12]. Human evaluation results in Table 2 demonstrate that our model achieves the best overall preference and prompt alignment among all methods, and AnyV2V (I2VGen-XL) is the most preferred method. We conjecture that the gain is coming from our compatibility with state-of-the-art image editing models.
We also employ automatic evaluation metrics on our edited video of the human evaluation datasets. Following previous works [5, 2], our automatic evaluation employs the CLIP [41] model to assess both text alignment and temporal consistency. For text alignment, we calculate the CLIP-Score, specifically by determining the average cosine similarity between the CLIP text embeddings derived from the editing prompt and the CLIP image embeddings across all frames. For temporal consistency, we evaluate the average cosine similarity between the CLIP image embeddings of every pair of consecutive frames. These two metrics are referred to as CLIP-Text and CLIP-Image, respectively. Our automatic evaluations in Table 2 demonstrate that our model is competitive in prompt-based editing compared to baseline methods.
@@
For novel tasks (2), (3) and (4), we evaluate the performance of three I2V generation models using human evaluations and show the results in Table 5. As these tasks require reference images instead of text prompts, we focus on evaluating the reference alignment and overall preference of the edited videos. According to the results, we observe that AnyV2V (I2VGen-XL) is the best model across all tasks, underscoring its robustness and versatility in handling diverse video editing tasks. AnyV2V (SEINE) and AnyV2V (ConsistI2V) show varied performance across tasks. AnyV2V (SEINE) performs good reference alignment in reference-based style transfer and identity manipulation, but falls short in subject-driven editing with lower scores. On the other hand, AnyV2V (ConsistI2V) shines in subject-driven editing, achieving second-best results in both reference alignment and overall preference. Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing [28], we also report the image editing success rate in Table 5 to clarify that our method relies on a good image frame edit.
B.3 Qualitative Results
@@
By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing. In our experiment, we used InstructPix2Pix [4] for the first frame edit. Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video. Comparing our work with the three baseline models TokenFlow [16], FLATTEN [12], and Tune-A-Video [53], the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt. The color tone and object shapes are also tilted. It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes. The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modeling.
@@
Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST [15] to obtain the edited frame. This level of controllability offers artists the unprecedented opportunity to use their art as a reference for video editing, opening new avenues for creative expression. As demonstrated in Figure 4, our method captures the distinctive style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately, while such an edit is often hard to perform using existing text-guided video editing methods.
@@
In our experiment, we employed a subject-driven image editing model AnyDoor [9] for the first frame editing. AnyDoor allows replacing any object in the target image with the subject from only one reference image. We observe from Figure 4 that AnyV2V produces highly motion-consistent videos when performing subject-driven object swapping. In the first example, AnyV2V successfully replaces the cat with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. In the second example, the car is replaced by our desired car while maintaining the rotation angle in the edited video.
@@
By integrating the identity-preserved image personalization model InstantID [49] with ControlNet [58], this approach enables the replacement of an individual’s identity to create an initial frame. Our AnyV2V framework then processes this initial frame to produce an edited video, swapping the person’s identity as showcased in Figure 4. To the best of our knowledge, our work is the first to provide such flexibility in the video editing models. Note that the InstantID with ControlNet method will alter the background due to its model property. It is possible to leverage other identity-preserved image personalization models and apply them to AnyV2V to preserve the background.
B.4 Human Evaluation
B.4.1 Dataset
Our human evaluation dataset contains a total of 89 samples that have been collected from https://www.pexels.com. For prompt-based editing, we employed InstructPix2Pix [4] to compose the examples. Topics include swapping objects, adding objects, and removing objects. For subject-driven editing, we employed AnyDoor [9] to replace objects with reference subjects. For Neural Style Transfer, we employed NST [15] to compose the examples. For identity manipulation, we employed InstantID [49] to compose the examples. See Table 4 for the statistic.
@@ B.4.1 Dataset
Prompt-based Editing 45 InstructPix2Pix [4]
Reference-based Style Transfer 20 NST [15]
Identity Manipulation 13 InstantID [49]
Subject-driven Editing 11 AnyDoor [9]
@@ B.4.2 Interface
Appendix C Preliminary
C.1 Image-to-Video (I2V) Generation Models
In this work, we focus on leveraging latent diffusion-based [43] I2V generation models for video editing. Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$. The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (3)
C.2 DDIM Inversion
The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM [44] sampling algorithm. The reverse process of DDIM sampling, known as DDIM inversion [35, 13], allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
C.3 Plug-and-Play (PnP) Diffusion Features
Tumanyan et al. [46] proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g. legs or torso of a human body) during the image generation process.
@@ C.3 Plug-and-Play (PnP) Diffusion Features
Image Editing Method NST [15] AnyDoor [9] InstantID [49]
@@
Appendix D Discussion
D.1 Limitations
Inaccurate Edit from Image Editing Models. As our method relies on an initial frame edit, the image editing models are used. However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [28]. For example, in the subject-driven video editing task, we found that AnyDoor [9] requires several tries to get a good editing result. Efforts are required in manually picking a good edited frame. We expect that in the future better image editing models will minimize such effort.
@@ D.1 Limitations
D.2 License of Assets
For image editing models, InstructPix2Pix [4] inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion. Neural Style Transfer [15] is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License. InstantID [49] is under Apache License 2.0. AnyDoor [9] is under the MIT License.
For baselines, Tune-A-Video [53] is under Apache License 2.0, TokenFlow [16] is under MIT License andFLATTEN [12] is under Apache License 2.0.

### v4

@@ 5.3 Evaluation Results
As shown in Figure 3 and 4, AnyV2V can perform the following video editing tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation. we compare AnyV2V against three baseline models Tune-A-Video (Wu et al., 2023b), TokenFlow (Geyer et al., 2023) and FLATTEN (Cong et al., 2023) for (1) prompt-based editing. Since there exists no publicly available baseline method for task (2) (3) (4), we evaluate the performance of three I2V generation models under AnyV2V. We included a more comprehensive evaluation in Appendix B.
Prompt-based Editing
@@
Unlike the baseline methods (Wu et al., 2023b; Geyer et al., 2023; Cong et al., 2023) that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged. Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix (Brooks et al., 2023), AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue. Additionally, it maintains the original video’s background and fidelity, whereas, in comparison, baseline methods often alter the color tone and shape of objects, as illustrated in Figure 3. Also, for motion tasks such as adding snowing weather, I2V models from AnyV2V provide inherent support for animating the snowing while the baseline methods would result in flickering. For quantitative evaluations, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting, and also compute the metrics CLIP-Text for text alignment and CLIP-Image for temporal consistency. In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model (Radford et al., 2021) and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames. Table 2 shows that AnyV2V generally achieves high text-alignment and temporal consistency, while AnyV2V with I2VGen-XL backbone is the most preferred method because it does not edit the video precisely.
Style Transfer, Subject-Driven Editing and Identity Manipulation
@@
For these novel tasks, we stress the alignment with the reference image instead of the text prompt. As shown in Figure 4, we can observe that for task (2), AnyV2V can capture one particular style that is tailor-made, and even if the style is not learned by the text encoder. In the examples, AnyV2V captures the style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately. For task (3), AnyV2V can replace the subject in the video with other subjects even if the new subject differs from the original subject. In the examples, a cat is replaced with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. A car is replaced by our desired car while the wheel is still spinning in the edited video. For task (4), AnyV2V can swap a person’s identity to anyone. We report both human evaluation results and find that AnyV2V with I2VGen-XL backbone is the most preferred method in terms of reference alignment and overall performance, plotted in Table 5, which is in Appendix B.
I2V Backbones
@@
We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively. It has a good generalization ability to produce consistent motions in the video with high visual quality. AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its training data, thus harming the visual quality. AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
@@
Figure 5: AnyV2V can edit video length beyond the training frame while maintaining motion consistency. The first row is the source video frames while the second rows are the edited. The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix (Brooks et al., 2023).
Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models (Chen et al., 2023d; Ren et al., 2024; Zhang et al., 2023c) are mostly trained on video data that contains only 16 frames. To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d). However, we find that such an experiment setup cannot maintain semantic consistency in our case. As many works in extending video generation exploit the initial latent to generate longer video (Qiu et al., 2023; Wu et al., 2023c), we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output. Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
@@ A.1 U-Net Layers for Feature Injection
To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL (Zhang et al., 2023c), ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d). Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e. average attention weights for all other tokens attending to the current token). The results are shown in Figure 7.
@@ A.2 Ablation Analysis on Feature Injection Thresholds
Appendix B Evaluation Detail
B.1 Quantitative Evaluations
@@
For (1) prompt-based editing, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting. We compare AnyV2V against three baseline models: Tune-A-Video (Wu et al., 2023b), TokenFlow (Geyer et al., 2023) and FLATTEN (Cong et al., 2023). Human evaluation results in Table 2 demonstrate that our model achieves the best overall preference and prompt alignment among all methods, and AnyV2V (I2VGen-XL) is the most preferred method. We conjecture that the gain is coming from our compatibility with state-of-the-art image editing models.
We also employ automatic evaluation metrics on our edited video of the human evaluation datasets. Following previous works (Ceylan et al., 2023; Bai et al., 2024), our automatic evaluation employs the CLIP (Radford et al., 2021) model to assess both text alignment and temporal consistency. For text alignment, we calculate the CLIP-Score, specifically by determining the average cosine similarity between the CLIP text embeddings derived from the editing prompt and the CLIP image embeddings across all frames. For temporal consistency, we evaluate the average cosine similarity between the CLIP image embeddings of every pair of consecutive frames. These two metrics are referred to as CLIP-Text and CLIP-Image, respectively. Our automatic evaluations in Table 2 demonstrate that our model is competitive in prompt-based editing compared to baseline methods.
@@
For novel tasks (2), (3) and (4), we evaluate the performance of three I2V generation models using human evaluations and show the results in Table 5. As these tasks require reference images instead of text prompts, we focus on evaluating the reference alignment and overall preference of the edited videos. According to the results, we observe that AnyV2V (I2VGen-XL) is the best model across all tasks, underscoring its robustness and versatility in handling diverse video editing tasks. AnyV2V (SEINE) and AnyV2V (ConsistI2V) show varied performance across tasks. AnyV2V (SEINE) performs good reference alignment in reference-based style transfer and identity manipulation, but falls short in subject-driven editing with lower scores. On the other hand, AnyV2V (ConsistI2V) shines in subject-driven editing, achieving second-best results in both reference alignment and overall preference. Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing (Ku et al., 2024), we also report the image editing success rate in Table 5 to clarify that our method relies on a good image frame edit.
B.2 Qualitative Results
@@
By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing. In our experiment, we used InstructPix2Pix (Brooks et al., 2023) for the first frame edit. Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video. Comparing our work with the three baseline models TokenFlow (Geyer et al., 2023), FLATTEN (Cong et al., 2023), and Tune-A-Video (Wu et al., 2023b), the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt. The color tone and object shapes are also tilted. It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes. The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modeling.
@@
Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST (Gatys et al., 2015) to obtain the edited frame. This level of controllability offers artists the unprecedented opportunity to use their art as a reference for video editing, opening new avenues for creative expression. As demonstrated in Figure 4, our method captures the distinctive style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately, while such an edit is often hard to perform using existing text-guided video editing methods.
@@
In our experiment, we employed a subject-driven image editing model AnyDoor (Chen et al., 2023c) for the first frame editing. AnyDoor allows replacing any object in the target image with the subject from only one reference image. We observe from Figure 4 that AnyV2V produces highly motion-consistent videos when performing subject-driven object swapping. In the first example, AnyV2V successfully replaces the cat with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. In the second example, the car is replaced by our desired car while maintaining the rotation angle in the edited video.
@@
By integrating the identity-preserved image personalization model InstantID (Wang et al., 2024b) with ControlNet (Zhang et al., 2023b), this approach enables the replacement of an individual’s identity to create an initial frame. Our AnyV2V framework then processes this initial frame to produce an edited video, swapping the person’s identity as showcased in Figure 4. To the best of our knowledge, our work is the first to provide such flexibility in the video editing models. Note that the InstantID with ControlNet method will alter the background due to its model property. It is possible to leverage other identity-preserved image personalization models and apply them to AnyV2V to preserve the background.
B.3 Human Evaluation
B.3.1 Dataset
Our human evaluation dataset contains a total of 89 samples that have been collected from https://www.pexels.com. For prompt-based editing, we employed InstructPix2Pix (Brooks et al., 2023) to compose the examples. Topics include swapping objects, adding objects, and removing objects. For subject-driven editing, we employed AnyDoor (Chen et al., 2023c) to replace objects with reference subjects. For Neural Style Transfer, we employed NST (Gatys et al., 2015) to compose the examples. For identity manipulation, we employed InstantID (Wang et al., 2024b) to compose the examples. See Table 4 for the statistic.
@@ B.3.1 Dataset
Prompt-based Editing 45 InstructPix2Pix
Reference-based Style Transfer 20 NST
Identity Manipulation 13 InstantID
Subject-driven Editing 11 AnyDoor
@@ B.3.2 Interface
Image Editing Method NST AnyDoor InstantID
@@
Appendix C Discussion
C.1 Limitations
Inaccurate Edit from Image Editing Models. As our method relies on an initial frame edit, the image editing models are used. However, the current state-of-the-art models are not mature enough to perform accurate edits consistently (Ku et al., 2024). For example, in the subject-driven video editing task, we found that AnyDoor (Chen et al., 2023c) requires several tries to get a good editing result. Efforts are required in manually picking a good edited frame. We expect that in the future better image editing models will minimize such effort.
@@ C.1 Limitations
C.2 License of Assets
For image editing models, InstructPix2Pix (Brooks et al., 2023) inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion. Neural Style Transfer (Gatys et al., 2015) is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License. InstantID (Wang et al., 2024b) is under Apache License 2.0. AnyDoor (Chen et al., 2023c) is under the MIT License.
For baselines, Tune-A-Video (Wu et al., 2023b) is under Apache License 2.0, TokenFlow (Geyer et al., 2023) is under MIT License andFLATTEN (Cong et al., 2023) is under Apache License 2.0.